# Cones of Noether–Lefschetz divisors and moduli spaces of hyperkähler manifolds

### v2

@@ Abstract.
[…]
@@ 1. Introduction
The moduli space $\mathcal{F}_{2d}$ arises as a quotient $\mathcal{D}_{\Lambda_{2d}}/\widetilde{{\rm{O}}}^{+}\left(\Lambda_{2d}\right)$, where $\mathcal{D}_{\Lambda_{2d}}$ is a Type IV symmetric domain associated to the even lattice of signature $(2,n)$ given by
and $\widetilde{{\rm{O}}}^{+}\left(\Lambda_{2d}\right)$, the group of orientation-preserving isomorphisms of $\Lambda$ which act trivially on the discriminant group $D(\Lambda_{2d})=\Lambda_{2d}^{\vee}\big{/}\Lambda_{2d}$. Heegner divisors generalize Noether–Lefschetz divisors to an arbitrary orthogonal modular varieties $\mathcal{D}_{\Lambda}/\Gamma$ associated to an even lattice $\Lambda$ of signature $(2,n)$, where $\Gamma$ is a finite index subgroup of $\widetilde{{\rm{O}}}^{+}\left(\Lambda\right)$, by viewing Noether–Lefschetz divisors as images of hyperplane arrangements in $\mathcal{D}_{\Lambda}$ under the modular projection $\pi:\mathcal{D}_{\Lambda}\longrightarrow\mathcal{D}_{\Lambda}\big{/}\Gamma$.
Bergeron–Li–Millson–Moeglin [BLMM17] and Bruinier–Zuffetti [BZ24] proved a generalization of Maulik–Pandharipande’s conjecture, showing that when $n\geq 3$ and $\Lambda$ splits off two copies of the hyperbolic plane, the Picard group with rational coefficients ${\rm Pic}_{\mathbb{Q}}(X)$ of any orthogonal modular variety $X=\mathcal{D}_{\Lambda}/\widetilde{{\rm{O}}}^{+}\left(\Lambda\right)$ is generated by Heegner divisors. The rank of ${\rm Pic}_{\mathbb{Q}}(\mathcal{D}_{\Lambda}/\Gamma)$ was computed by Bruinier in [Bru02b].
For any orthogonal modular variety $X=\mathcal{D}_{\Lambda}/\Gamma$ as above, the NL cone ${\rm{Eff}}^{NL}\left(X\right)\subset{\rm{Pic}}_{\mathbb{Q}}\left(X\right)$ is the convex cone of effective $\mathbb{Q}$-linear combinations of irreducible components of Heegner divisors (known as primitive Heegner divisors) on $X$. The NL cone contains the subcone ${\rm{Eff}}^{H}\left(\mathcal{D}_{\Lambda}/\Gamma\right)$ generated by the (non-primitive) Heegner divisors on $\mathcal{D}_{\Lambda}/\Gamma$. After tensoring with $\mathbb{R}$, this NL cone ${\rm{Eff}}^{NL}\left(X\right)$ forms a natural subcone of the cone of pseudo-effective divisors $\overline{{\rm Eff}}(X)$.
@@
Bruinier–Möller [BM19] answered the first question affirmatively, showing that for any orthogonal modular variety $X=\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda\right)$ with $n\geq 3$ splitting off two copies of the hyperbolic plane, the cone ${\rm{Eff}}^{NL}\left(X\right)$ is always polyhedral.
In this paper, we tackle Question (2) for $X=\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda\right)$ under the same assumptions. We consider the $\mathbb{Q}$-vector space $S_{k,\Lambda}$ of vector-valued cusp forms of weight $k=1+\frac{n}{2}$ with respect to the Weil representation [Bor98] and the coefficient extraction functionals in $S_{k,\Lambda}^{\vee}$
$c_{m,\mu}:S_{k,\Lambda}\longrightarrow\mathbb{Q},\quad\sum a_{m,\mu}q^{m}\mathfrak{e}_{\mu}\mapsto a_{m,\mu}.$
Let $b\geq\lceil k/12\rceil$ be an integer such that the set of $c_{m,\mu}$ with $0<m\leq b$ and $\mu\in D(\Lambda)$ generates $S_{k,\Lambda}^{\vee}$. Then, we consider the weakly holomorphic modular form
@@
where $\Delta(\tau)$ is the scalar-valued discriminant modular form and $E_{(2-k)+12b,\Lambda(-1)}$ is the Eisenstein series of weight $(2-k)+12b$ associated to $\Lambda(-1)$ (see Equation (5)).
@@
Let $\Lambda$ be an even lattice of signature $(2,n)$ with $n\geq 4$ splitting off two copies of the hyperbolic plane and $X=\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}(\Lambda)$ its modular variety. Fixing $b$ as above, there are explicit bounds $\Xi$ and $\Omega$, depending on $k$, the discriminant of $\Lambda$, and the $\alpha_{m,\mu}$ with $-b\leq m\leq 0$ in (1), such that
@@
Theorem 1.1 together with its implementation in Sage package [heegner_cones] enables the description of ${\rm{Eff}}^{NL}\left(X\right)$ in terms of generators for any such $\Lambda$ (see Section 1.1 below).
We remark that the bound $\Xi$ of Theorem 3.4 in fact allows for the description of the cone in $\left({\rm Mod}^{\circ}_{k,\Lambda}\right)^{\vee}$ generated by coefficient extraction functionals $c_{m,\mu}\colon{\rm Mod}^{\circ}_{k,\Lambda}\rightarrow\mathbb{Q}$ under the weaker assumption that $\Lambda$ splits off only one copy of the hyperbolic plane. Here ${\rm Mod}^{\circ}_{k,\Lambda}=\mathbb{Q}E_{k,\Lambda}\oplus S_{k,\Lambda}$ is the $\mathbb{Q}$-vector space of almost cups forms. The assumption that $\Lambda$ splits off two copies of the hyperbolic plane is needed to convert the above result into a description of ${\rm{Eff}}^{NL}\left(X\right)$ via the results of [BLMM17, BZ24].
@@ 1.1. Applications to moduli
[…]
In some cases, one can use the position of the canonical class $K_{X}$ with respect to the NL cone to show that $X$ has negative Kodaira dimension. This occurs when $K_{X}$ lies on the opposite side from the NL cone of the hyperplane in ${\rm Pic}_{\mathbb{Q}}(X)$ of divisors with degree 00 with respect to the Hodge class $\lambda$. We formalize this condition numerically in terms of the Eisenstein series (see the more general Proposition 4.2) in order to give the following criterion for uniruledness.
@@
then the orthogonal modular variety $X=\mathcal{D}_{\Lambda}\big{/}\widetilde{{\rm{O}}}^{+}(\Lambda)$ is uniruled.
@@ 1.2. Uniruledness results
Mukai in a celebrated series of papers [Muk88, Muk92, Muk06, Muk10, Muk16] constructed unirational parameterizations of $\mathcal{F}_{2d}$ for low-degrees. This has been recently improved by Farkas–Verra in [FV18, FV21]. The first examples of higher dimensional projective hyperkähler varieties were exhibited in [Bea83], they are generalized Kummer varieties and Hilbert schemes of points on K3 surfaces. When considered polarized they deform in $4$ and $20$ dimensional moduli spaces respectively.
[…]
@@
    when $\gamma=2$ for $t\leq 10$ and $t=12$ with $d=4t-1$,
@@
    when $\gamma=2$ for $t\leq 9$ and $t=11,13$ with $d=4t-2$.
@@
    when $\gamma=2$ for $t\leq 11$ and $t=13,15,17,19$, where $n=4t-2$.
@@ Acknowledgements
This paper benefited from helpful discussions and correspondence with the following people who we gratefully acknowledge: Daniele Agostini, Emma Brakkee, Jan Hendrik Bruinier, Yagna Dutta, Gabi Farkas, Paul Kiefer, Giovanni Mongardi, Gregory Sankaran, Preston Wake, and Riccardo Zuffetti.
@@ 2. Preliminaries
$\left\{[Z]\in\mathbb{P}\left(\Lambda_{\mathbb{C}}\right)\left|\langle Z,Z\rangle=0,\langle Z,\overline{Z}\rangle>0\right.\right\}.$
Further, we denote by $\Gamma$ a finite index subgroup of the group ${\rm{O}}^{+}\left(\Lambda\right)$ of automorphisms of $\Lambda$ fixing the component $\mathcal{D}_{\Lambda}$. The quotient of $\mathcal{D}_{\Lambda}$ by $\Gamma$ called an orthogonal modular variety. It is a quasi-projective variety [BB66] that for various choices of lattice $\Lambda$ and arithmetic groups $\Gamma$ gives a partial compactification of a coarse moduli space of polarized varieties. The first case of interest in this paper is when
@@
$\mathcal{F}_{2d}=\mathcal{D}_{\Lambda_{2d}}\big{/}\widetilde{{\rm{O}}}^{+}\left(\Lambda_{2d}\right)$
@@
As mentioned in the introduction, a very general point $(S,H)\in\mathcal{F}_{2d}$ has Picard group ${\rm{Pic}}(S)=\mathbb{Z}H$, and a large source of geometric divisors comes from imposing the condition that the Picard rank jumps. These are Noether–Lefschetz divisors. There are different characterizations of these divisors: by keeping track of a rank two lattice embedding $L\hookrightarrow{\rm{Pic}}(S)$, by imposing the existence of an extra class $\beta\in{\rm{Pic}}(S)$ with fixed intersections $(\beta^{2},\beta\cdot H)=(2h-2,a)$, and by looking at images of hyperplanes in $\mathcal{D}_{\Lambda_{2d}}$ via the quotient map
$\pi_{2d}:\mathcal{D}_{\Lambda_{2d}}\longrightarrow\mathcal{F}_{2d}.$
@@ 2.1. Heegner and NL divisors
$D_{v}=v^{\perp}\cap\mathcal{D}_{\Lambda}=\left\{[Z]\in\mathcal{D}_{\Lambda}\left|\langle Z,v\rangle=0\right.\right\}.$
@@
Under the given assumption that $\Gamma\subset\widetilde{O}^{+}\left(\Lambda\right)$, all the components of $H_{m,\mu}$ have multiplicity two if $\mu=-\mu$ in $\Lambda^{\vee}\big{/}\Lambda$ and all have multiplicity one otherwise. Further, the line bundle $\mathcal{O}(-1)$ on $\mathcal{D}_{\Lambda}\subset\mathbb{P}\left(\Lambda_{\mathbb{C}}\right)$ admits a natural $\Gamma$-action and descends to a $\mathbb{Q}$-line bundle $\lambda$ called the Hodge bundle. One declares $H_{0,0}=-\lambda$.
In the K3 case $\mathcal{F}_{2d}=\mathcal{D}_{\Lambda_{2d}}\big{/}\widetilde{{\rm{O}}}^{+}\left(\Lambda_{2d}\right)$, Noether-Lefschetz divisors are often described as the reduced divisor obtained by taking the closure of the locus
@@
of points $(S,H)$ for which there exists a class $\beta\in{\rm{Pic}}(S)$ not proportional to $H$ with $\beta^{2}=2h-2$ and $\beta\cdot H=a$. In this case [MP13]*Lemma 3, if $d$ does not divides $a$:
@@
[BLMM17, Theorem 1.5], [BZ24, Remark 3.13, Corollary 3.18] Let $\Lambda$ be an even lattice of signature $(2,n)$ with $n\geq 3$ splitting off two copies of the hyperbolic plane. Then the rational Picard group of $\mathcal{D}_{\Lambda}\big{/}\widetilde{{\rm{O}}}^{+}\left(\Lambda\right)$ is generated by Heegner divisors:
${\rm{Pic}}_{\mathbb{Q}}^{H}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{{\rm{O}}}^{+}\left(\Lambda\right)\right)={\rm{Pic}}_{\mathbb{Q}}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{{\rm{O}}}^{+}\left(\Lambda\right)\right).$
Note that the above theorem in particular implies that irreducible components of $H_{m,\mu}$ must be linear combinations of other Heegner divisors. This relation is explicit and follows from Eichler’s criterion [GHS09]*Proposition 3.3, [Son23]*Proposition 2.15: if $\Lambda$ splits off two copies of the hyperbolic lattice $U$, then the $\widetilde{\rm{SO}}^{+}\left(\Lambda\right)$-orbit of a primitive element $v\in\Lambda^{\vee}$ is determined by $Q(v)=m$ and $v+\Lambda\in\Lambda^{\vee}\big{/}\Lambda$. This leads to the following definition (see [Pet15, BM19]). The primitive Heegner divisor $P_{\Delta,\delta}$ is the image via the $\Gamma$-quotient map $\pi:\mathcal{D}_{\Lambda}\longrightarrow\mathcal{D}_{\Gamma}\big{/}\Gamma$ of the cycle
@@
When $\Lambda$ splits off two copies of $U$, and $\Gamma=\widetilde{\rm{O}}^{+}\left(\Lambda\right)$, the divisor $P_{\Delta,\delta}$ is irreducible and reduced when $\delta\neq-\delta$ in $D(\Lambda)$ and otherwise has multiplicity two. The relation between Heegner and primitive Heegner divisors [BM19]*Equations (17) and (18) is:
@@
where the $\mu(\cdot)$ in the second equation stands for the Möbius function. Here $r^{2}\mid m$ means exactly that there is a class $\delta\in D(\Lambda)$ such that $m/r^{2}\in Q(\delta)+\mathbb{Z}$.
As stated in the introduction, our main object of study is the NL cone ${\rm{Eff}}^{NL}\left(\mathcal{D}_{\Lambda}\big{/}\Gamma\right)$ generated by primitive Heegner divisors, or equivalently irreducible components of Noether-Lefschetz divisors.

### v3

@@ Abstract.
We give a general formula for generators of the NL cone on an orthogonal modular variety. This is the cone of effective divisors linearly equivalent to an effective linear combination of irreducible components of Noether-Lefschetz divisors. We apply this to describe, in terms of minimal generators, the NL cone of various moduli spaces of geometric origin such as those of polarized K3 surfaces, cubic fourfolds, and hyperkähler manifolds. Additionally, we establish uniruledness for many moduli spaces of primitively polarized hyperkähler manifolds of $\rm{OG6}$} and ${\rm{Kum}}_{n}$-type. Finally, in analogy with the case of K3 surfaces of degree $2$, we show that any family of polarized ${\rm{Kum}}_{2}$-type hyperkähler manifolds with divisibility $2$ and polarization degree $2$ over a projective base is isotrivial.
@@ 1. Introduction
The moduli space $\mathcal{F}_{2d}$ arises as a quotient $\mathcal{D}/\widetilde{{\rm{O}}}^{+}\left(\Lambda_{2d}\right)$, where $\mathcal{D}$ is the Type IV symmetric domain associated with ${\rm O}(2,n)$, the lattice $\Lambda_{2d}$ is the even lattice of signature $(2,n)$ given by
$\Lambda_{2d}=U^{\oplus 2}\oplus E_{8}(-1)^{\oplus 2}\oplus\mathbb{Z}\ell,\;\;\hbox{with}\;\;\langle\ell,\ell\rangle=-2d,$
and $\widetilde{{\rm{O}}}^{+}\left(\Lambda_{2d}\right)$ is the group of orientation-preserving isomorphisms of $\Lambda_{2d}$ which act trivially on the discriminant group $D(\Lambda_{2d})=\Lambda_{2d}^{\vee}\big/\Lambda_{2d}$. For an arbitrary orthogonal modular variety $\mathcal{D}/\Gamma$, where $\Gamma$ is a finite index subgroup of $\widetilde{{\rm{O}}}^{+}\left(\Lambda\right)$ and $\Lambda$ an even lattice of signature $(2,n)$, Noether–Lefschetz divisors generalize to Heegner divisors, which are images of hyperplane arrangements in $\mathcal{D}$ under the modular projection $\pi:\mathcal{D}\longrightarrow\mathcal{D}\big/\Gamma$.
Bergeron–Li–Millson–Moeglin [BLMM17] and Bruinier–Zuffetti [BZ24] proved a generalization of Maulik–Pandharipande’s conjecture, showing that when $n\geq 3$ and $\Lambda$ splits off two copies of the hyperbolic plane, the Picard group with rational coefficients ${\rm Pic}_{\mathbb{Q}}(X)$ of any orthogonal modular variety $X=\mathcal{D}/\widetilde{{\rm{O}}}^{+}\left(\Lambda\right)$ is generated by Heegner divisors. The rank of ${\rm Pic}_{\mathbb{Q}}(\mathcal{D}/\Gamma)$ was computed by Bruinier in [Bru02b].
For any orthogonal modular variety $X=\mathcal{D}/\Gamma$ as above, the NL cone ${\rm{Eff}}^{NL}\left(X\right)\subset{\rm{Pic}}_{\mathbb{Q}}\left(X\right)$ is the convex cone of effective $\mathbb{Q}$-linear combinations of irreducible components of Heegner divisors (known as primitive Heegner divisors) on $X$. The NL cone contains the subcone ${\rm{Eff}}^{H}\left(\mathcal{D}/\Gamma\right)$ generated by the (non-primitive) Heegner divisors on $\mathcal{D}/\Gamma$. After tensoring with $\mathbb{R}$, this NL cone ${\rm{Eff}}^{NL}\left(X\right)$ forms a natural subcone of the cone of pseudo-effective divisors $\overline{{\rm Eff}}(X)$.
@@
Bruinier–Möller [BM19] answered the first question affirmatively, showing that for any orthogonal modular variety $X=\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda\right)$ with $n\geq 3$ splitting off two copies of the hyperbolic plane, the cone ${\rm{Eff}}^{NL}\left(X\right)$ is always polyhedral.
In this paper, we tackle Question (2) for $X=\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda\right)$ under the same assumptions. We consider the $\mathbb{Q}$-vector space $S_{k,\Lambda}$ of vector-valued cusp forms of weight $k=1+\frac{n}{2}$ with respect to the Weil representation [Bor98] and the coefficient extraction functionals in $S_{k,\Lambda}^{\vee}$
$c_{m,\mu}:S_{k,\Lambda}\longrightarrow\mathbb{Q},\quad\sum a_{m,\mu}q^{m}\mathfrak{e}_{\mu}\mapsto a_{m,\mu},$
where $\{\mathfrak{e}_{\mu}\mid\mu\in D(\Lambda)\}$ are the standard generators of the group algebra $\mathbb{C}\left[D(\Lambda)\right]$. Let $b\geq\lceil k/12\rceil$ be an integer such that the set of $c_{m,\mu}$ with $0<m\leq b$ and $\mu\in D(\Lambda)$ generates $S_{k,\Lambda}^{\vee}$. Then, we consider the weakly holomorphic modular form
@@
where $\Delta(\tau)$ is the scalar-valued discriminant modular form and $E_{(2-k)+12b,\Lambda(-1)}$ is the Eisenstein series of weight (2k)+12b(2-k)+12b associated to $\Lambda(-1)$ (see Equation (5)).
@@
Let $\Lambda$ be an even lattice of signature $(2,n)$ with $n\geq 4$ splitting off two copies of the hyperbolic plane and $X=\mathcal{D}\big/\widetilde{\rm{O}}^{+}(\Lambda)$ its modular variety. Fixing $b$ as above, there are explicit bounds $\Xi$ and $\Omega$, depending on $k$, the discriminant of $\Lambda$, and the $\alpha_{m,\mu}$ with $b\leq m\leq 0$ in (1), such that
@@
Theorem 1.1 together with its implementation in a Sage package [heegner_cones] enables the description of ${\rm{Eff}}^{NL}\left(X\right)$ in terms of generators for any such $\Lambda$ (see Section 1.1 below).
We remark that the bound $\Xi$ of Theorem 3.4 in fact allows for the description of the cone in $\left({\rm Mod}^{\circ}_{k,\Lambda}\right)^{\vee}$ generated by coefficient extraction functionals $c_{m,\mu}\colon{\rm Mod}^{\circ}_{k,\Lambda}\rightarrow\mathbb{Q}$ under the weaker assumption that $\Lambda$ splits off only one copy of the hyperbolic plane. Here ${\rm Mod}^{\circ}_{k,\Lambda}=\mathbb{Q}E_{k,\Lambda}\oplus S_{k,\Lambda}$ is the $\mathbb{Q}$-vector space of almost cusp forms. The assumption that $\Lambda$ splits off two copies of the hyperbolic plane is needed to convert the above result into a description of ${\rm{Eff}}^{NL}\left(X\right)$ via the results of [BLMM17, BZ24].
@@ 1.1. Applications to moduli
We then focus on cases where the quotient $X=\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda\right)$ arises as a partial compactification of a coarse moduli space of polarized K3 surfaces, hyperkähler manifolds, or cubic fourfolds. We give explicit formulas for ${\rm{Eff}}^{NL}\left(X\right)$ in terms of generating rays for low-degree polarizations: see Table LABEL:sec6:table:NLK3 for the case of (quasi)-polarized K3 surfaces and Tables 2 and 3 for the case of hyperkähler fourfolds of ${\rm K3}^{[2]}$-type. We remark that the orthogonal modular variety $X=\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda\right)$ partially compactifying the moduli space of smooth cubic fourfolds is the same as that partially compactifying the moduli space of polarized hyperkähler manifolds of ${\rm K3}^{[2]}$-type with polarization of divisibility $2$ and degree $6$ and thus the description of ${\rm{Eff}}^{NL}\left(X\right)$ for cubic fourfolds is already contained in Table 3. In the case of $\mathcal{F}_{2d}$, the calculations in Table LABEL:sec6:table:NLK3 confirm (aside from one additional necessary generator in the case $d=13$) the predictions in [Pet15] who computed, for $d\leq 18$, the cone generated by the set of $8d$ primitive Heegner divisors $P_{\Delta,\delta}$, for $\delta\in D(\Lambda),\Delta\in Q(\delta)+s$ with $s=0,1,2,3$, and conjectured that this cone coincides with the one generated by all of them, that is, with ${\rm{Eff}}^{NL}\left(\mathcal{F}_{2d}\right)$.
In some cases, one can use the position of the canonical class $K_{X}$ with respect to the NL cone to show that $X$ has negative Kodaira dimension. This occurs when $K_{X}$ lies on the opposite side from the NL cone of the hyperplane in ${\rm Pic}_{\mathbb{Q}}(X)$ of divisors with degree 0 with respect to the Hodge class $\lambda$. We formalize this condition numerically in terms of the Eisenstein series (see the more general Proposition 4.2) in order to give the following criterion for uniruledness.
@@
then the orthogonal modular variety $X=\mathcal{D}\big/\widetilde{{\rm{O}}}^{+}(\Lambda)$ is uniruled.
@@ 1.2. Uniruledness results
Mukai in a celebrated series of papers [Muk88, Muk92, Muk06, Muk10, Muk16] constructed unirational parameterizations of $\mathcal{F}_{2d}$ for low degrees. This has been recently improved by Farkas–Verra in [FV18, FV21]. The first examples of higher dimensional projective hyperkähler varieties were exhibited in [Bea83]: they are generalized Kummer varieties and Hilbert schemes of points on K3 surfaces. Polarized varieties of these two types deform in $4$ and $20$ dimensional moduli spaces respectively.
The problem of exhibiting a projective realization of a generic such object is intimately related with rationality properties of the corresponding moduli space. Although constructions of ${\rm Kum}_{n}$ and ${\rm{K3}}^{[n]}$-type hyperkähler varieties were exhibited at the same time, unirational parameterizations are available only for some moduli spaces of hyperkähler varieties of ${\rm{K3}}^{[n]}$-type, see [BD85, OG06, IR01, IR07, DV10, BLM+21]. Constructing unirational parameterizations in low degree for moduli spaces of other types of hyperkähler varieties (e.g. generalized Kummer, ${\rm{OG}}6$, and ${\rm{OG}}10$ types) presents a challenge where, as far as we know, not a single explicit construction is known. Here we consider the simpler problem of establishing uniruledness and focus on the generalized Kummer and ${\rm{OG}}6$-type cases (our methods do not yield new results in the case of ${\rm{OG}}10$-type).
@@
    when $\gamma=2$ for $t\leq 10$ and $t=12$ with d=4t1d=4t-1,
@@
    when $\gamma=2$ for $t\leq 9$ and $t=11,13$ with d=4t2d=4t-2.
@@
    when $\gamma=2$ for $t\leq 11$ and $t=13,15,17,19$, where n=4t2n=4t-2.
@@ Acknowledgements
This paper benefited from helpful discussions and correspondence with the following people who we gratefully acknowledge: Daniele Agostini, Emma Brakkee, Jan Hendrik Bruinier, Yagna Dutta, Gabi Farkas, Paul Kiefer, Giovanni Mongardi, Gregory Sankaran, Preston Wake, and Riccardo Zuffetti. We would also like to thank the anonymous referee for helpful comments and corrections.
@@ 2. Preliminaries
Let $\Lambda$ be an even lattice of signature $(2,n)$ with bilinear form given by $\langle\cdot,\cdot\rangle$. The bilinear form extends to $\Lambda_{\mathbb{C}}$ and we let $\mathcal{D}$ denote one of the two components of
@@
Further, we denote by $\Gamma$ a finite index subgroup of the group ${\rm{O}}^{+}\left(\Lambda\right)$ of automorphisms of $\Lambda$ fixing the component $\mathcal{D}$. The quotient of $\mathcal{D}$ by $\Gamma$ is called an orthogonal modular variety. It is a quasi-projective variety [BB66] which for some choices of lattice $\Lambda$ and arithmetic groups $\Gamma$ gives a partial compactification of a coarse moduli space of polarized varieties. The first case of interest in this paper is when
@@
$\mathcal{F}_{2d}=\mathcal{D}\big/\widetilde{{\rm{O}}}^{+}\left(\Lambda_{2d}\right)$
@@
As mentioned in the introduction, a very general point $(S,H)\in\mathcal{F}_{2d}$ has Picard group ${\rm{Pic}}(S)=\mathbb{Z}H$, and a large source of geometric divisors comes from imposing the condition that the Picard rank jumps. These are Noether–Lefschetz divisors. There are different characterizations of these divisors: by keeping track of a rank two lattice embedding $L\hookrightarrow{\rm{Pic}}(S)$, by imposing the existence of an extra class $\beta\in{\rm{Pic}}(S)$ with fixed intersections $(\beta^{2},\beta\cdot H)=(2h-2,a)$, and by looking at images of hyperplanes in $\mathcal{D}$ via the quotient map
$\pi_{2d}:\mathcal{D}\longrightarrow\mathcal{F}_{2d}.$
@@ 2.1. Heegner and NL divisors
[…]
@@
Under the given assumption that $\Gamma\subset\widetilde{\rm{O}}^{+}\left(\Lambda\right)$, all the components of $H_{m,\mu}$ have multiplicity two if $\mu=-\mu$ in $\Lambda^{\vee}\big/\Lambda$ and all have multiplicity one otherwise. Further, the line bundle $\mathcal{O}(-1)$ on $\mathcal{D}\subset\mathbb{P}\left(\Lambda_{\mathbb{C}}\right)$ admits a natural $\Gamma$-action and descends to a $\mathbb{Q}$-line bundle $\lambda$ called the Hodge bundle. One declares $H_{0,0}=-\lambda$.
In the K3 case $\mathcal{F}_{2d}=\mathcal{D}\big/\widetilde{{\rm{O}}}^{+}\left(\Lambda_{2d}\right)$, Noether-Lefschetz divisors are often described as the reduced divisor obtained by taking the closure of the locus
@@
of points $(S,H)$ for which there exists a class $\beta\in{\rm{Pic}}(S)$ not proportional to $H$ with $\beta^{2}=2h-2$ and $\beta\cdot H=a$. In this case [MP13]*Lemma 3, if $d$ does not divide $a$:
@@
[BLMM17, Theorem 1.5], [BZ24, Remark 3.13, Corollary 3.18] Let $\Lambda$ be an even lattice of signature $(2,n)$ with $n\geq 3$ splitting off two copies of the hyperbolic plane. Then the rational Picard group of $\mathcal{D}\big/\widetilde{{\rm{O}}}^{+}\left(\Lambda\right)$ is generated by Heegner divisors:
${\rm{Pic}}_{\mathbb{Q}}^{H}\left(\mathcal{D}\big/\widetilde{{\rm{O}}}^{+}\left(\Lambda\right)\right)={\rm{Pic}}_{\mathbb{Q}}\left(\mathcal{D}\big/\widetilde{{\rm{O}}}^{+}\left(\Lambda\right)\right).$
Note that the above theorem in particular implies that irreducible components of $H_{m,\mu}$ must be linear combinations of other Heegner divisors. This relation is explicit and follows from Eichler’s criterion [GHS09]*Proposition 3.3, [Son23]*Proposition 2.15: if $\Lambda$ splits off two copies of the hyperbolic lattice $U$, then the $\widetilde{\rm{SO}}^{+}\left(\Lambda\right)$-orbit of a primitive element $v\in\Lambda^{\vee}$ is determined by $Q(v)=m$ and $v+\Lambda\in\Lambda^{\vee}\big/\Lambda$. This leads to the following definition (see [Pet15, BM19]). The primitive Heegner divisor $P_{\Delta,\delta}$ is the image via the $\Gamma$-quotient map $\pi:\mathcal{D}\longrightarrow\mathcal{D}\big/\Gamma$ of the cycle
@@
Under the assumption that $\Lambda$ splits off two copies of $U$ and $\Gamma=\widetilde{\rm{O}}^{+}\left(\Lambda\right)$, the divisor $P_{\Delta,\delta}$ is always irreducible. Moreover, in this situation, the divisor $P_{\Delta,\delta}$ is reduced when $\delta\neq-\delta$ in $D(\Lambda)$ and has multiplicity two otherwise. The relation between Heegner and primitive Heegner divisors [BM19]*Equations (17) and (18) is:
@@
where $r^{2}\mid m$ denotes that there is a class $\delta\in D(\Lambda)$ such that $m/r^{2}\in Q(\delta)+\mathbb{Z}$ and the $\mu(\cdot)$ in the second equation stands for the Möbius function.
As stated in the introduction, our main object of study is the NL cone ${\rm{Eff}}^{NL}\left(\mathcal{D}\big/\Gamma\right)$ generated by primitive Heegner divisors, or equivalently irreducible components of Noether-Lefschetz divisors.